# Deep spatial context: when attention-based models meet spatial regression

### 18 Jan 2024 v1

@@ 1 Introduction
The behavior and reasoning of Deep Learning (DL) models are often hard to understand, especially for people without experience in artificial intelligence (AI). This is the source of a low level of trust in such models, especially in critical domains, e.g. healthcare. Physicians would like to know if DL models take similar issues into account when making a diagnosis as they do. The need to understand the models was one of the motivations behind the eXplainable AI (XAI) raise Guidotti et al. (2018). However, as pointed out in Longo et al. (2023), most existing XAI methods do not operate on the same level of abstraction as people interested in the understanding of systems. For instance, gradient-based methods Selvaraju et al. (2017); Sundararajan et al. (2017) can highlight particular pixels within images, e.g. constituting face, but this is not clear whether it is the shape, texture, color, etc. that matters. That is why we need to operate on semantic human-understandable concepts when explaining models. It is best if the concepts are specified by experts from the domain which the system is designed for. The issue was pointed out in the manifesto of XAI 2.0 (human-centered explanations) Longo et al. (2023).
@@
The topic of context within the input data is mainly considered in time series and natural language processing, where it is understood as the number of data pieces surrounding the analyzed data sample. The fact that the size of the context matters is outlined in Tworkowski et al. (2023). The concept of context in the case of images is not widely covered. It can be perceived as additional tabular information attached to the images. However, in this work, we focus on spatial context within images which we define as the fact that the location of particular elements of the image and their relative positions are significant in a decision-making process. For instance, the fact of how much the areas with tumor are spread is an important factor when assessing the disease stage. Therefore, it translates to the risk stratification and treatment Dee et al. (2023). The other illustrative example to give the intuition behind the idea of spatial context is the scenario where there are two people in an image and one has a knife. Then, there is a big difference in whether the person with a knife is heading another person or going in the opposite direction. Therefore, the relative position of the two entities matters.
@@
The common approach to the classification of histopathological images in the form of Whole Slide Images is attention-based models. One of such solutions is CLAM Lu et al. (2021). In this case, the image of huge resolution is first split into patches that are the input to the feature extractor. The features are then aggregated in the pooling module, where the assigned attention scores serve as coefficients in a weighted sum to form an overall representation of the entire original image. Such a representation is an input to the final classifier. In such attention-based models, the information about the patches’ location within the whole image is not explicitly preserved, as no positional encoding is applied. This can be troublesome in the case of histopathological images that have non-standardized sizes.
@@
We propose a ‘Deep spatial context’ method (DSCon) (Figure 1) which is in line with the notion of XAI 2.0 Longo et al. (2023) as it investigates the DL model from the point of view of the semantic concept of spatial context. The research was inspired by histopathologists, although the method is general and can be applied to other domains. In the solution, we use spatial regression models to investigate the DL model’s internals.
@@ 2 Related works
Many works in the XAI field focus on highlighting the parts of the input image that lead to the largest neuron activations and therefore influence the final model decisions the most. In the early works, the contribution of single input pixels (saliency maps Simonyan et al. (2013)) and superpixel (groups of pixels) Ribeiro et al. (2016) were mainly developed. The recent works focus more on high-level, complex semantic concepts Kim et al. (2017); Achtibat et al. (2023); Sun et al. (2023). There are also neural networks that are explainable-by-design providing trainable prototypes Chen et al. (2019); Rymarczyk et al. (2023). In all these solutions, the process is that the XAI methods return some explanations that require humans to name them as human-understandable, semantic concepts i.e. bill, eye.
@@
In the papers from the field of digital pathology, the analysis of the model behavior is often limited to information about the global metrics like accuracy and AUC accompanied by few heatmaps with patches’ importance in the overall tissue classification. There are only a few attempts to explain the models in more detail (based on prototypes Rymarczyk et al. (2023) and an idea similar to global centroids Xie et al. (2020)). To the best of our knowledge, there are no works that analyse the models from the spatial context perspective.
The attention mechanism is widely used in different forms in DL systems, however, it was pointed out that the attention scores themselves are not valid model explanations Jain and Wallace (2019). There are recent works where the attention mechanism is used in another manner to explain models. The idea is to insert the external attention mechanism into the already trained convolutional model to train it to derive class activation maps being a form of model explanation Gkartzonika et al. (2023); Ntrougkas et al. (2022).
However, there have been only a few attempts to investigate the way the attention mechanism works from a more theoretic point of view. The self-attention can be considered as a marginal likelihood over possible network structures Singh and Buckley (2023), via a dynamical system perspective Huang et al. (2023) or through the lens of exchangeability Zhang et al. (2022). To the best of our knowledge, we are the first to apply spatial statistics for the analysis of attention-based models and Deep Learning models, in general. However, there are already XAI methods that inherit from different mathematical subdisciplines - e.g. SHAP values Lundberg and Lee (2017) are inspired by game theory.
@@ 3.1 Preliminaries - spatial regression
The two common spatial regression classes of models are (1) spatial lag model ($y=\rho(Wy)+W\beta+\epsilon$) i.e in a form of spatial two-stage least squares model Anselin (1988) (denoted as lag method for brevity), (2) spatial error model ($y=X\beta+u$ where $u=\lambda(Wu)+\epsilon$) i.e in a form of generalized method of moments for a spatial error model Kelejian and Prucha (1999, 1998) (denoted as error method). Moreover, the OLS model can be applied with manually lagged features based on the neighboring patches - SLX ($y=X\beta+WX\theta+\epsilon$) (referred as Wx). Lagging means that we use a transformation to incorporate the information from the neighboring areas (more in the Appendix). This model is actually ‘pseudo-spatial’; however, we will not use this distinction in the following sections.
@@ 3.2 Deep Spatial Context
The next step is to check whether the spatial regression is applicable. The common indicator is that the residuals of OLS form spatial clusters. It can be formally verified by performing Global Moran’s I test Moran (1948) which checks if there is a global spatial autocorrelation in the input data (see Appendix). At this stage, the spatial weights matrix ($W$) has to be defined i.e. using $k$-nearest neighbors. Later, it is checked whether the $p$-value returned in the test is smaller than the defined significance level. If so, the application of spatial regression is justifiable. Using the same set of features as in OLS, spatial regression models (Wx, lag, error) are run and the returned performance metrics are saved.
@@ 4.1 Deep Learning models
In all illustrative experiments, we used the Camelyon16 Breast Cancer dataset Ehteshami Bejnordi et al. (2017) containing 399 WSIs in total. Inspired by the Rashomon’s quartet  Biecek et al. (2023), where several models of similar performance proved to reason about the data in different ways, we decided to test the proposed ‘Deep spatial context’ method on a range of different attention-based models achieving similar classification performance to verify whether, in this case, the differences in capturing the spatial context are significant.
In the experiments, we use CLAM implementation of the attention-based model Lu et al. (2021). We replaced the baseline Resnet50 feature extractor with different transformer-based backbones to verify how the extracted features vary and later impact the learned attention scores which are the core input to the DSCon. We applied the Swin model Liu et al. (2021) also in the second version Liu et al. (2022) which are transformer models that resemble the convolutional models due to a hierarchical structure. Furthermore, we used Vision Transformer (ViT) Dosovitskiy et al. (2021). Models from Swin and ViT families differ in terms of the scope of attention - in the first case, it is local (window-based), whereas, in the second it is global (self-attention).
@@ 4.2 Training
The initial parameters of the grid search were inspired by the ones used in Jia et al. (2022). Later, they were modified based on the intermediate results. After an extensive search over many hyperparameters, we chose those that for every model variant (with different backbones) resulted in a decent performance similar to the baseline with Resnet50 backbone (denoted as resnetCLAM) (details in the Appendix). In resnetCLAM, the parameters claimed in the original paper were applied. To conclude, the models used in the analysis varied in terms of architecture only by the choice of the backbone. For brevity, the whole trained attention-based models (CLAM) are named after the feature extractor used. The classification performance of the models is summarized in Figure 2. The differences in performance among the models are not statistically significant.
Figure 2: Critical distance plot [^2] The code used for the generation of the plot is from Ismail Fawaz et al. (2019).of models’ accuracy of classification of normal tissue vs. tumor lesion. For brevity: s denotes Swin, sv2 - Swin v2 and w - window. In the labels, the mean accuracy (percentage) and the standard deviation over 5 trained models during the cross-validation are provided. The fact that the horizontal line below the ranking axis connects all the models means that there is no significant difference in their performance based on the Wilcoxon-Holm method used to detect pairwise significance.
@@ 5.6 Local inspection of the features-attention scores relationship
Spatial statistics allow not only to analyze data from the global point of view but also from the local one. One of the methods is to use a local indicator of spatial autocorrelation (LISA) Anselin (1995) which is formally similar to the Global Moran’s I statistic but the summation over all regions is removed. We provide a visualization of one of ‘the most spatial’ images (‘test_001.tif’) from the local perspective using LISA (Figure 7).
@@ A.1 Training details
After an extensive search over a grid of hyperparameters, the following parameters were chosen for the final experiments on models with transformer-based backbones: learning_rate=0.005, weight_decay_rate=0, dropout_rate=0.25, bag_weight=0.7. In the first layer of the classifier, there were 512 neurons, whereas 128 in the second. Following the setup in Jia et al. [2022], the SGD optimizer was used.

### 10 Mar 2024 v2

@@ 1 Introduction
The behavior and reasoning of Deep Learning (DL) models are often hard to understand, especially for people without experience in artificial intelligence (AI). This is the source of a low level of trust in such models, especially in critical domains, e.g. healthcare. Physicians would like to know if DL models take similar issues into account when making a diagnosis as they do. The need to understand the models was one of the motivations behind the eXplainable AI (XAI) raise [1]. However, as pointed out in [2], most existing XAI methods do not operate on the same level of abstraction as people interested in the understanding of systems. For instance, gradient-based methods [3, 4] can highlight particular pixels within images, e.g. constituting face, but this is not clear whether it is the shape, texture, color, etc. that matters. That is why we need to operate on semantic human-understandable concepts when explaining models. It is best if the concepts are specified by experts from the domain which the system is designed for. The issue was pointed out in the manifesto of XAI 2.0 (human-centered explanations) [2].
@@
The topic of context within the input data is mainly considered in time series and natural language processing, where it is understood as the number of data pieces surrounding the analyzed data sample. The fact that the size of the context matters is outlined in [5]. The concept of context in the case of images is not widely covered. It can be perceived as additional tabular information attached to the images. However, in this work, we focus on spatial context within images which we define as the fact that the location of particular elements of the image and their relative positions are significant in a decision-making process. For instance, the fact of how much the areas with tumor are spread is an important factor when assessing the disease stage. Therefore, it translates to the risk stratification and treatment [6]. The other illustrative example to give the intuition behind the idea of spatial context is the scenario where there are two people in an image and one has a knife. Then, there is a big difference in whether the person with a knife is heading another person or going in the opposite direction. Therefore, the relative position of the two entities matters.
@@
The common approach to the classification of histopathological images in the form of Whole Slide Images is attention-based models. One of such solutions is CLAM [7]. In this case, the image of huge resolution is first split into patches that are the input to the feature extractor. The features are then aggregated in the pooling module, where the assigned attention scores serve as coefficients in a weighted sum to form an overall representation of the entire original image. Such a representation is an input to the final classifier. In such attention-based models, the information about the patches’ location within the whole image is not explicitly preserved, as no positional encoding is applied. This can be troublesome in the case of histopathological images that have non-standardized sizes.
@@
We propose a ‘Deep spatial context’ method (DSCon) (Figure 1) which is in line with the notion of XAI 2.0 [2] as it investigates the DL model from the point of view of the semantic concept of spatial context. The research was inspired by histopathologists, although the method is general and can be applied to other domains. In the solution, we use spatial regression models to investigate the DL model’s internals.
@@ 2 Related works
Many works in the XAI field focus on highlighting the parts of the input image that lead to the largest neuron activations and therefore influence the final model decisions the most. In the early works, the contribution of single input pixels (saliency maps [8]) and superpixel (groups of pixels) [9] were mainly developed. The recent works focus more on high-level, complex semantic concepts [10, 11, 12]. There are also neural networks that are explainable-by-design providing trainable prototypes [13, 14]. In all these solutions, the process is that the XAI methods return some explanations that require humans to name them as human-understandable, semantic concepts i.e. beak, eye.
@@
In the papers from the field of digital pathology, the analysis of the model behavior is often limited to information about the global metrics like accuracy and AUC accompanied by few heatmaps with patches’ importance in the overall tissue classification. There are only a few attempts to explain the models in more detail (based on prototypes [14] and an idea similar to global centroids [15]). To the best of our knowledge, there are no works that analyse the models from the spatial context perspective.
The attention mechanism is widely used in different forms in DL systems, however, it was pointed out that the attention scores themselves are not valid model explanations [16]. There are recent works where the attention mechanism is used in another manner to explain models. The idea is to insert the external attention mechanism into the already trained convolutional model to train it to derive class activation maps being a form of model explanation [17, 18].
However, there have been only a few attempts to investigate the way the attention mechanism works from a more theoretic point of view. The self-attention can be considered as a marginal likelihood over possible network structures [19], via a dynamical system perspective [20] or through the lens of exchangeability [21]. To the best of our knowledge, we are the first to apply spatial statistics for the analysis of attention-based models and Deep Learning models, in general. However, there are already XAI methods that inherit from different mathematical subdisciplines - e.g. SHAP values [22] are inspired by game theory.
@@ 3.1 Preliminaries - spatial regression
The two common spatial regression classes of models are (1) spatial lag model ($y=\rho(Wy)+W\beta+\epsilon$) i.e in a form of spatial two-stage least squares model [23] (denoted as lag method for brevity), (2) spatial error model ($y=X\beta+u$ where $u=\lambda(Wu)+\epsilon$) i.e in a form of generalized method of moments for a spatial error model [24, 25] (denoted as error method). Moreover, the OLS model can be applied with manually lagged features based on the neighboring patches - SLX ($y=X\beta+WX\theta+\epsilon$) (referred as Wx). Lagging means that we use a transformation to incorporate the information from the neighboring areas (more in the Appendix). This model is actually ‘pseudo-spatial’; however, we will not use this distinction in the following sections.
@@ 3.2 Deep Spatial Context
The next step is to check whether the spatial regression is applicable. The common indicator is that the residuals of OLS form spatial clusters. It can be formally verified by performing Global Moran’s I test [26] which checks if there is a global spatial autocorrelation in the input data (see Appendix). At this stage, the spatial weights matrix ($W$) has to be defined i.e. using $k$-nearest neighbors. Later, it is checked whether the $p$-value returned in the test is smaller than the defined significance level. If so, the application of spatial regression is justifiable. Using the same set of features as in OLS, spatial regression models (Wx, lag, error) are run and the returned performance metrics are saved.
@@ 4.1 Deep Learning models
In all illustrative experiments, we used the Camelyon16 Breast Cancer dataset [27] containing 399 WSIs in total. Inspired by the Rashomon’s quartet  [28], where several models of similar performance proved to reason about the data in different ways, we decided to test the proposed ‘Deep spatial context’ method on a range of different attention-based models achieving similar classification performance to verify whether, in this case, the differences in capturing the spatial context are significant.
In the experiments, we use CLAM implementation of the attention-based model [7]. We replaced the baseline Resnet50 feature extractor with different transformer-based backbones to verify how the extracted features vary and later impact the learned attention scores which are the core input to the DSCon. We applied the Swin model [29] also in the second version [30] which are transformer models that resemble the convolutional models due to a hierarchical structure. Furthermore, we used Vision Transformer (ViT) [31]. Models from Swin and ViT families differ in terms of the scope of attention - in the first case, it is local (window-based), whereas, in the second it is global (self-attention).
@@ 4.2 Training
The initial parameters of the grid search were inspired by the ones used in [32]. Later, they were modified based on the intermediate results. After an extensive search over many hyperparameters, we chose those that for every model variant (with different backbones) resulted in a decent performance similar to the baseline with Resnet50 backbone (denoted as resnetCLAM) (details in the Appendix). In resnetCLAM, the parameters claimed in the original paper were applied. To conclude, the models used in the analysis varied in terms of architecture only by the choice of the backbone. For brevity, the whole trained attention-based models (CLAM) are named after the feature extractor used. The classification performance of the models is summarized in Figure 2. The differences in performance among the models are not statistically significant.
Figure 2: Critical distance plot [^2] The code used for the generation of the plot is from [33].of models’ accuracy of classification of normal tissue vs. tumor lesion. For brevity: s denotes Swin, sv2 - Swin v2 and w - window. In the labels, the mean accuracy (percentage) and the standard deviation over 5 trained models during the cross-validation are provided. The fact that the horizontal line below the ranking axis connects all the models means that there is no significant difference in their performance based on the Wilcoxon-Holm method used to detect pairwise significance.
@@ 5.6 Local inspection of the features-attention scores relationship
Spatial statistics allow not only to analyze data from the global point of view but also from the local one. One of the methods is to use a local indicator of spatial autocorrelation (LISA) [34] which is formally similar to the Global Moran’s I statistic but the summation over all regions is removed. We provide a visualization of one of ‘the most spatial’ images (‘test_001.tif’) from the local perspective using LISA (Figure 7).
@@ A.1 Training details
After an extensive search over a grid of hyperparameters, the following parameters were chosen for the final experiments on models with transformer-based backbones: learning_rate=0.005, weight_decay_rate=0, dropout_rate=0.25, bag_weight=0.7. In the first layer of the classifier, there were 512 neurons, whereas 128 in the second. Following the setup in [32], the SGD optimizer was used.